# references.bib RefAV: Towards Planning-Centric Scenario Mining

### v2

@@ 4 Experiments
In this section, we briefly describe our evaluation metrics (Sec. 4.1), provide an empirical analysis of our baselines (Sec. 4.2), and ablate the impact of different LLMs on RefProg’s spatio-temporal scenario mining accuracy (Sec. 4.3).
4.1 Metrics
We evaluate scenario mining methods using HOTA-Temporal and HOTA-Track (variants of HOTA [luiten2021hota]) to measure referential tracking performance and balanced accuracy to measure information retrieval accuracy. HOTA is a unified metric that explicitly balances detection accuracy, association, and localization, making it better aligned with human visual evaluations of tracking performance. HOTA-Temporal extends the standard HOTA metric by only considering the referred timestamps within a track as true positives. HOTA-Track is similar to HOTA-Temporal, but it does not penalize methods for incorrectly predicting the start and end of a referred action. For example, given the prompt car turning right, HOTA-Temporal only considers the timestamps where the car is turning right in the full referred track as true positives, whereas HOTA-Track considers all timestamps in the referred track as true positives. We also include HOTA results to contextualize standard tracking performance with the referential tracking performance of all benchmarked trackers. Additionally, we use timestamp balanced accuracy as a timestamp-level classification metric and log balanced accuracy as a log-level classification metric to evaluate how well methods can identify which timestamps and logs contain objects that match the referential prompt, respectively.
For each prompt, we categorize all objects into three groups: referred objects, which are the primary objects specified by the prompt; related objects, which are objects that interact with the referred object; and other objects, which are neither referred to nor relevant to the prompt. We compute both HOTA-Temporal and HOTA-Track metrics exclusively for the referred object class in the main paper. We report the performance on related objects and other objects in Appendix C.
@@ 4.1 Metrics
| Filtering by Referred Class (RefBlind) | |||||
@@
| Referential Track Inflation (RefInflate) | |||||
| GroundingDINO [liu2024grounding] | - | 4.0 | 6.0 | 49.7 | 49.3 |
| Referential Track Classification (RefClassify) | |||||
@@
| Referential Tracking by Program Synthesis (RefProg) | |||||
@@ 4.2 Empirical Analysis of Results
Table 2 presents several simple baselines for addressing spatio-temporal scenario mining. We evaluate ground-truth oracles for each zero-shot referential tracker to understand how tracking accuracy and language grounding independently influence final performance. First, evaluating all ground truth tracks without considering language prompts yields a HOTA-Temporal of 13.3%. Further, evaluating RefBlind with ground truth tracks improves referential tracking accuracy by 8%. However, RefBlind with LE3DE2E tracks results in a slight performance drop. Next, we find that RefInflate performs the worst out of all four baselines, as inflating 2D detections to 3D is challenging. In contrast, RefClassify performs considerably better. Interestingly, the RefClassify oracle outperforms the RefBlind oracle, suggesting that CLIP-based filtering can capture richer semantics than class names alone. However, RefClassify underperforms RefBlind when using predicted tracks (e.g., LE3DE2E scores 17.2 vs. 19.2 HOTA-Temporal). Finally, our proposed RefProg baseline significantly outperforms all other zero-shot methods. Notably, RefProg with LE3DE2E tracks achieves a 32.8% improvement in HOTA-Temporal over RefClassify with Le3DE2E tracks.
Analysis of Failure Cases. We find that all three baselines have unique failure modes. First, we note that RefInflate often predicts bounding boxes even in cases where no referred object exists. We attribute this to GroundingDINO’s [liu2024grounding] referential grounding pre-training, which biases the model towards always predicting a bounding box since each training image always has at least one true positive grouth-truth box. Next, we find that RefClassify performs poorly because the CLIP image embedding for each track lacks the context required to understand multi-agent interactions. Specifically, the 2D projected image only includes the tracked object, and lacks temporal context. Lastly, we find that RefProg fails in cases where the API listing cannot generate a valid program for a given prompt (e.g. prompts involving weather and lighting conditions). Although one can always add new atomic actions, this strategy is not scalable.
@@ 4.3 Ablation Study
Impact of LLMs on Code Generation. We evaluate the impact of using different LLMs on code generation quality in RefProg. We use Le3DE2E’s tracks for all experiments in this ablation. Somewhat unsurprisingly, we find that Claude 3.7 Sonnet performs the best (cf. Table 3), achieving a HOTA-Temporal of 50.1. We posit that this is because it has been explicitly tuned for code generation and instruction following. Gemini 2.0 Flash performs considerably worse despite also being tuned for coding. Interestingly, we find that Gemini and Qwen 2.5 Instruct struggle to generate valid programs, and instead try to generate full programs.
4.4 CVPR 2025 Scenario Mining Challenge
We hosted a challenge at CVPR 2025 to encourage broad community involvement in addressing spatio-temporal scenario mining. Our competition recieved submissions from eight teams (some submissions are private) at the close of our competition on June 8th, 2025 AOE. Notably, four teams beat our best baseline. We present the current top three teams at the bottom of Table 2. We summarize the contributions of the top three teams in Appendix F and include a link to full technical reports and code here.
4.5 Limitations and Future Work
@@
Noisy Ground Truth Perception Labels. The quality of our scenario mining dataset is limited by the quality of the ground truth perception labels in AV2. For example, jittery tracks can lead to poor motion classification over short horizons. We address this issue by significantly post-processing and manually verifying all generated scenarios to mitigate label noise.
Limited Diversity. Although AV2 includes many interesting scenarios (cf. Fig. 4), it is still relatively small compared to industry-scale datasets. Moreover, the AV2 sensor dataset was primarily created to train and benchmark 3D perception models, and is not intended for scenario mining. As a result, it does not include many useful annotations like traffic light state, weather information, or lighting annotations. Although our dataset is of similar size to prior work like nuPrompt, future datasets should be explicitly curated to address scenario mining.
Runtime Complexity. RefProg requires running offline perception algorithms and executing LLM-based programs, both of which are prohibitively slow. Although these components are trivially parallelizable, we posit that our approach is not sufficiently efficient for real-world applications. Future work should also consider the runtime and scalability of their method beyond logs.
@@ 5 Conclusion
6 Acknowledgments
This work was supported in part by funding from Bosch Research and the NSF GRFP (Grant No. DGE2140739). We would also like to thank Ishan Khatri for his feedback on early drafts
@@ Appendix A Implementation Details
We present additional implementation details to reproduce our baseline experiments below. Our code is available on GitHub. We primarily consider two-stage baselines that first track all objects, then classify them into referred objects, related objects, or other objects. We repurpose 3D trackers from winning teams in the Argoverse 2 End-to-End Forecasting Challenge, and include a link to the base tracker outputs on Google Drive. All models are trained to output tracks for 26 annotated categories within Argoverse 2. We include an example prediction below:
We append a track for the ego vehicle with a consistent track_id, confidence of 1.00, class name of EGO_VEHICLE, yaw of 0 radians, size (<length, width height> in meters) of <4.877, 2, 1.473>, and translation (<x, y, z> in meters) of <1.422, 0, 0.25>. This translation represents the offset from the centroid of the ego-vehicle to the ego-vehicle reference coordinate located near the center of its rear axle.
@@
Filtering by Referred Class (RefBlind). RefBlind filters all tracks except those corresponding to the referred class. To get the referred class, we prompt Claude 3.7 Sonnet as follows:
Categories: {AV2 Categories} Please select the one category that most corresponds to the object of focus in the description:{Natural Language Description}. As an example, for the description "vehicle turning at intersection with nearby pedestrians" your output would be VEHICLE. For the description "ego vehicle near construction barrel" your output would be EGO_VEHICLE. Your only output should be the category name, in all capital letters and including underscores if necessary.
@@
Referential Track Classification (RefClassify). RefClassify filters tracks by computing prompt-track CLIP similarity over time. At each timestep, we project the eight vertices of each object’s bounding box onto the image plane of all seven cameras. We then cull any objects that do not partially fall within the frustum of each camera. For each projected bounding box, if three or more vertices fall within the image plane, we take a crop using the minimum and maximum uv image coordinates. We pad this crop by 30 pixels on each side and feed the resulting image into a CLIP ViT-L/14 model to obtain CLIP image embeddings. We also use CLIP ViT-L/14 to compute text embeddings for all prompts in the dataset. Next, we calculate cosine similarity scores between each track’s image crops and the text prompt. Since individual similarity scores tend to be noisy, we follow ReferGPT [chamiti2025refergpt] in applying a dynamic threshold and majority voting strategy to identify the referred object tracks. Since the dataset includes both positive and negative examples for each prompt, we group similarity scores by prompt and threshold the top 10% of scores. If more than half of the scores of a track fall within the top 10% threshold, we label the entire track as a referred object. To reduce the number of length one tracks (i.e., tracks spanning only one timestamp), we apply a similarity score modifier: .
Referential Track Inflation (RefInflate). RefInflate leverages off-the-shelf foundational models to enable zero-shot 3D referential tracking. We implement a GroundingSAM-like pipeline [liu2024grounding, kirillov2023segment], where GroundingDino generates bounding boxes based on a given prompt, and SAM2 segments objects within those bounding boxes. We extend this process to predict 3D bounding boxes using multi-view video and LiDAR data [khurana2024shelf]. For each camera, we prompt GroundingDino with the input image and referential prompt, and extract bounding boxes that we then use to prompt SAM2 for object segmentation. We perform all operations at the native image resolution.
Next, we project all LIDAR points within each camera’s frustum onto the corresponding image plane and retain only the points that fall within the SAM2 segmentation mask. From the retained points, we compute the medoid. We then apply DBSCAN to cluster these LiDAR points, using parameters and . We further filter the points to include only those in the same cluster as the medoid. If DBSCAN identifies the medoid as noise, we instead select the largest cluster. We then draw a tight-fitting oriented bounding box around the filtered set of LiDAR points. To simplify orientation estimation, we align the front of the bounding box with the front of the ego vehicle.
To associate bounding boxes across time, we use a greedy tracker that applies Hungarian Matching to match boxes from timestep to tracks from using world-coordinate distance. We set the maximum matching distance to meters.
@@
Referential Tracking by Program Synthesis (RefProg). RefProg filters off-the-shelf tracks using Python programs synthesized by a large language model (LLM), without explicitly relying on any information from LiDAR or cameras. To reduce the time required to run these synthesized programs, we first filter the input tracks based on confidence. We compute each track’s confidence score by summing its confidence values across all timestamps. For the classes REGULAR_VEHICLE, PEDESTRIAN, BOLLARD, CONSTRUCTION_CONE, and CONSTRUCTION_BARREL, we retain the top 200 tracks per class. For all other classes, we retain the top 100 tracks per category. We prompt the LLM with the following prompt:
@@
Please use the following functions to find instances of a referred object in an autonomous driving dataset. Be precise to the description, try to avoid returning false positives. API Listing: {API Listing} Categories: {AV2 Categories} Define a single scenario for the description:{Natural Language Description} Here is a list of examples: {Prediction Examples}. Only output code and comments as part of a Python block. Feel free to use a liberal amount of comments. Do not define any additional functions, or filepaths. Do not include imports. Assume the log_dir and output_dir variables are given. Wrap all code in one python block and do not provide alternatives. Output code even if the given functions are not expressive enough to find the scenario.
We include the API listing in Appendix G. We found this prompt consistently produces reasonable code across all tested LLMs. After running the synthesized programs, we post-process the output tracks in accordance with dataset practices. We remove all length one tracks as noise and discard object relationships where the spatial distance exceeds 50 meters.
@@
For the remaining tracks, we apply timestamp dilation to reduce flickering of the referred object labels. We symmetrically dilate time segments to a minimum of 1.5 seconds. For example, if the base tracker tracks an object with ID 1 from 1.0 to 15.0 seconds at 2 Hz, and the synthesized code marks it as referred at timestamps 4.5 and 5.0, we update the final output to label the track as referred from 4.0 to 5.5 seconds. Finally, we downsample all output tracks to 2 Hz for evaluation.
Appendix B Scenario Mining Annotation Tool

### v3

@@ 4 Experiments
In this section, we briefly describe our evaluation metrics, provide an empirical analysis of our baselines, ablate the impact of different LLMs on RefProg’s spatio-temporal scenario mining accuracy, and evaluate RefProg on nuPrompt to demonstrate the robustness of our approach.
@@
| Filter by Referred Class | |||||
@@
| Image-Embedding Similarity | |||||
@@
| ReferGPT [chamiti2025refergpt] | |||||
| Le3DE2E [chenle3de2e] | 74.4 | 20.2 | 30.8 | 57.0 | 57.1 |
| LLM APIs as a Black Box | |||||
| Le3DE2E [chenle3de2e] | 74.4 | 37.2 | 39.2 | 58.4 | 62.3 |
| Referential Tracking by Program Synthesis | |||||
@@
Metrics. We evaluate scenario mining methods using HOTA-Temporal and HOTA-Track (variants of HOTA [luiten2021hota]) to measure referential tracking performance and balanced accuracy to measure information retrieval accuracy. HOTA is a unified metric that explicitly balances detection accuracy, association, and localization, making it better aligned with human visual evaluations of tracking performance. HOTA-Temporal extends the standard HOTA metric by only considering the referred timestamps within a track as true positives. HOTA-Track is similar to HOTA-Temporal, but it does not penalize methods for incorrectly predicting the start and end of a referred action. For example, given the prompt car turning right, HOTA-Temporal only considers the timestamps where the car is turning right in the full referred track as true positives, whereas HOTA-Track considers all timestamps in the referred track as true positives. We also include HOTA results to contextualize standard tracking performance with the referential tracking performance of all benchmarked trackers. Additionally, we use timestamp balanced accuracy as a timestamp-level classification metric and log balanced accuracy as a log-level classification metric to evaluate how well methods can identify which timestamps and logs contain objects that match the referential prompt, respectively. Balanced accuracy is used over binary classification metrics such as F1 due to the imbalance of positive and negative prompts.
For each prompt, we categorize all objects into three groups: referred objects, which are the primary objects specified by the prompt; related objects, which are objects that interact with the referred object; and other objects, which are neither referred to nor relevant to the prompt. We compute both HOTA-Temporal and HOTA-Track metrics exclusively for the referred object class in the main paper. We report the performance on related objects and other objects in the Appendix E.
@@ 4.1 Empirical Analysis of Results
Table 2 presents several simple baselines for addressing spatio-temporal scenario mining. We evaluate each of the locally run zero-shot baselines using oracle ground-truth track annotations to understand how tracking accuracy and language grounding independently influence final performance. We do not evaluate ReferGPT and LLM APIs as a Black Box with the ground-truth track annotations to avoid dataset leakage. Evaluating every track as referred object without considering language prompts (first row) yields a HOTA-Temporal of 13.3% and HOTA-Track of 20.5%. Predicting the nearest object class that corresponds to the prompt with Filter by Referred Class improves referential tracking accuracy by 8%. However, Filter by Referred Class with LE3DE2E tracks results in a slight performance drop over ground truth tracks. This performance drop is perhaps lower than expected because Le3DE2E has exceptionally high tracking accuracy for REGULAR VEHICLE, which dominates the AV2 dataset. The Image-Embedding Similarity oracle outperforms the Filter by Referred Class oracle, suggesting that CLIP-based filtering can capture richer semantics than class names alone. However, Image-Embedding Similarity underperforms Filter by Referred Class when using predicted tracks (e.g., LE3DE2E scores 17.2 vs. 19.2 HOTA-Temporal). This likely stems from CLIP being much more sensitive to the predicted 3D location of the bounding box. ReferGPT slightly outperforms Filter by Referred Class and Image-Embedding Similarity across all reported metrics. ReferGPT, unlike the other methods, uses information about the current position and velocity of an object relative to the ego vehicle. Notably, the LLMs as a Black Box baseline significantly improves over prior approaches with a HOTA-Temporal of 37.2%. Finally, our proposed RefProg baseline outperforms all other zero-shot methods. Notably, RefProg with LE3DE2E tracks achieves a 13.8% improvement in HOTA-Temporal over LLMs as a Black Box with Le3DE2E tracks.
Scenario Mining Challenge. We hosted a challenge at CVPR 2025 to encourage broad community involvement in addressing spatio-temporal scenario mining. Our competition received submissions from eight teams. Notably, four teams beat our best baseline. We present the top three entries at the bottom of Table 2, summarize their contributions in Appendix H and include a link to full technical reports and code here. The best performing team beats our baseline by 3.3% achieving 53.4 HOTA-Temporal.
@@
Impact of LLMs on Code Generation. We evaluate the impact of using different LLMs on code generation quality in RefProg. We use Le3DE2E’s tracks and a consistent prompt for all experiments in this ablation. We find that Claude 3.7 Sonnet performs the best (cf. Table 3), achieving a HOTA-Temporal of 50.1. We posit that this is because it has been explicitly tuned for code generation and instruction following. Gemini 2.0 Flash performs considerably worse. Interestingly, we find that Gemini-2.5-Flash and Qwen-2.5-Instruct struggle to generate valid programs, and instead try to generate full programs using invalid import statements. Failure Rate indicates the percentage of generated programs that throw an exception while running.
Evaluating RefProg on nuPrompt. The atomic functions used in RefAV’s dataset generation and RefProg represent fundamental motion primitives that are generalizable across datasets. To this end, we evaluate RefProg on nuPrompt [wu2023nuprompt], a popular referential-tracking dataset based on nuScenes in Table 4. Notably, we find that our zero-shot compositional approach outperforms current state-of-the-art methods trained on domain-specific data. We make minimal adaptations to RefProg for the nuScenes dataset. We swap out the AV2 annotated class list for the nuScenes annotated class list and remove atomic functions that require access to an HD map. Importantly, we make zero modifications to our atomic action definitions, and expect that dataset-specific modifications can improve performance further.
| Method | Decoder | AMOTA | AMOTP | Recall | MOTA |
| CenterPoint Tracker [yin2021center] | DETR3D | 0.079 | 1.820 | 0.196 | 0.093 |
| CenterPoint Tracker [yin2021center] | PETR | 0.178 | 1.650 | 0.291 | 0.197 |
| DQTrack [li2023dqtrack] | DETR3D | 0.186 | 1.641 | 0.307 | 0.208 |
| DQTrack [li2023dqtrack] | Stereo | 0.198 | 1.625 | 0.309 | 0.214 |
| DQTrack [li2023dqtrack] | PETRv2 | 0.234 | 1.545 | 0.332 | 0.269 |
| ADA-Track [ding2024ada] | PETR | 0.249 | 1.538 | 0.353 | 0.270 |
| PromptTrack [wu2023nuprompt] | DETR3D | 0.202 | 1.615 | 0.310 | 0.222 |
| PromptTrack [wu2023nuprompt] | PETR | 0.259 | 1.513 | 0.366 | 0.280 |
| RefProg (Ours) | StreamPETR [wang2023exploring] | 0.321 | 1.238 | 0.504 | 0.329 |
Analysis of Failure Cases. While RefProg outperforms Image-Embedding Similarity and ReferGPT, each method has unique failure cases. We find that Image-Embedding Similarity performs poorly because the CLIP image embedding for each track lacks the context required to understand multi-agent interactions. Specifically, the 2D projected image only includes the tracked object, and does not contain information about past or future frames. ReferGPT similarly suffers from a lack of temporal context. Lastly, we find that RefProg fails in cases where the API listing is not expressive enough for a given prompt (e.g., prompts involving weather and lighting conditions). Although one can always add new atomic actions, this strategy is not scalable.
@@
Limitations and Future Work. The quality of our scenario mining dataset is limited by the quality of the ground truth 3D perception labels in AV2. For example, jittery tracks can lead to poor motion classification over short horizons. We address this issue by significantly post-processing and manually verifying all generated scenarios to mitigate label noise. Although AV2 includes many interesting scenarios (cf. Fig. 3), it is still relatively small compared to industry-scale datasets. Future datasets should be explicitly curated to address spatiotemporal scenario mining.
@@ Appendix A Implementation Details
We present additional implementation details to reproduce our baseline experiments below. We primarily consider two-stage baselines that first track all objects and then classify them into referred objects, related objects, or other objects. We repurpose 3D trackers from winning teams of the Argoverse 2 End-to-End Forecasting Challenge. All models are trained to output tracks for 26 annotated categories within Argoverse 2. We include an example prediction signature below:
We append a track for the ego vehicle with a consistent track_id, confidence of 1.00, class name of EGO_VEHICLE, yaw of 0 radians, size (<length, width height> in meters) of <4.877, 2, 1.473>, and translation (<x, y, z> in meters) of <1.422, 0, 0.25>. This translation represents the offset from the centroid of the ego-vehicle to the ego-vehicle reference coordinate near the rear axle.
@@
Filter by Referred Class. This baseline filters all tracks except those corresponding to the referred class. To get the referred class, we prompt Claude 3.7 Sonnet as follows:
Categories: {AV2 Categories} Please select the one category that most corresponds to the object of focus in the description:{Natural Language Description}. As an example, for the description “vehicle turning at intersection with nearby pedestrians” your output would be VEHICLE. For the description “ego vehicle near construction barrel” your output would be EGO_VEHICLE. Your only output should be the category name.
@@
Image Embedding Similarity. This baseline filters tracks by computing prompt-track CLIP similarity over time. At each timestep, we project the eight vertices of each object’s bounding box onto the image plane of all seven cameras. We then cull any objects that do not partially fall within the frustum of each camera. For each projected bounding box, if three or more vertices fall within the image plane, we take a crop using the minimum and maximum uv image coordinates. We pad this crop by 30 pixels on each side and feed the resulting image into a CLIP ViT-L/14 model to obtain CLIP image embeddings. We also use CLIP ViT-L/14 to compute text embeddings for all prompts in the dataset. Next, we calculate cosine similarity scores between each track’s image crops and the text prompt. Since individual similarity scores tend to be noisy, we follow ReferGPT [chamiti2025refergpt] in applying a dynamic threshold and majority voting strategy to identify the referred object tracks. Since the dataset includes both positive and negative examples for each prompt, we group similarity scores by prompt and threshold the top 10% of scores. If more than half of the scores of a track fall within the top 10% threshold, we label the entire track as a referred object. To reduce the number tracks of length one (i.e., tracks spanning only one timestamp), we apply a similarity score modifier: .
LLM APIs as a Black Box We use the OpenAI Responses API to determine whether particular scenarios occur within a data log. We prompt GPT-5 with three files: an HD map data, 6-DoF ego-to-city transformations, and object tracker predictions. We use gpt-5-2025-08-07 with medium reasoning effort and access to a code interpreter tool. We give instructions to save a CSV containing the referred object id’s and timestamps for each scenario. All 150 test data logs are run in parallel, with one request and one response for each data log. The saved CSVs are parsed into the RefAV format for evaluation as each response is delivered. We include the instructions given to GPT-5 below.
You are an expert data analyst. Your job is to identify if a given scenarios occurs within an autonomous driving log. A log contains files with an HD map, 6-DoF poses of the ego vehicle, and bounding box annotations from an object tracker. For each of the given scenario descriptions, output a CSV file containing the track_uuids and timestamps of the referred objects. Some logs may not contain any referred objects. Your csv file should have “track_uuid” and “timestamp_ns” columns. Your output csv files should have the naming convention scenario{i}.csv, depending on the order the descriptions are given. Make sure to save all of your CSVs! Identify if each of these scenarios occurs within the log. {scenario_descriptions}
ReferGPT. This baseline uses a VLM to caption the cropped images of each track at each timestep. Tracks are filtered by computing a text similarity score between the prompt and track captions. We take several steps to adapt ReferGPT to the RefAV dataset. At each timestep, we project the eight vertices of each predicted track’s bounding box onto the image plane of all seven cameras. We then cull any objects that do not partially fall within the frustum of each camera. For each projected bounding box, if three or more vertices fall within the image plane, we take a crop using the minimum and maximum uv image coordinates. If the predicted track is visible in multiple cameras, we use the 2D bounding box with the highest area. We pad this crop by 10 of the crop’s length on each side. For the ego-vehicle, the crop is the entire image from the front camera. Each of the per object-timestamp crops is also paired with the motion statistics. We compute a track ’s motion statistics at timestep as , where x and y are the position of the track in meters, is the yaw of the track in degrees, and are the velocities of the track in meters per second, is the yaw rate of the track in degrees per second. is the distance from the centroids of the ego-vehicle and track in meters. All values are with respect to a static ego-centric coordinate frame. We prompt gpt-5-mini to caption each of the images as follows:
Caption this image taken from an autonomous vehicle (referred to as the ego-vehicle). Our detection model classifies this object as {object_category}. Here are the coordinates of the object {object_motion}. This is in the format [x, y, yaw, x_velocity, y_velocity, angular_velocity, distance]. This is in a coordinate system where positive x is to the front of the ego vehicle, positive y is to the left of the ego-vehicle, and positive z is above the ego-vehicle. Position and velocity are measured in meters and meters per second. Yaw is measured in degrees counter-clockwise (colloquially “left”) from the x-axis. Angular velocity is in degrees per second. Distance is distance from the ego-vehicle. Please describe the appearance of the object and scene, if they are located relative to the road/ego-vehicle, how they are moving, and how they are interacting with any nearby objects (if any are visible). Treat values of x, y, x_velocity, and y_velocity near zero as zero. Feel free to add more relevant information in the captions.
We present an example caption below:
A motorcyclist slightly to the front-right of the ego-vehicle is stopped or creeping forward beside a compact SUV, straddling a scooter with feet down and hands on the handlebars. The rider is close to the SUV’s rear quarter, appearing stationary or moving very slowly while sharing the lane alongside other stopped traffic. The scene looks like an urban intersection or congested street with vehicles side-by-side and the rider positioned between them.
To determine if an object is referred, we compute a score . is the cosine similarity between siglip2-base-patch16-naflex text embeddings of the caption and prompt. We follow the original ReferGPT implementation and compute using Gestalt pattern matching to find the similarity between two strings. Finally, we determine that an entire track is referred if the majority the track’s caption scores are above .
We find that the SigLIPv2 text similarity does not effectively separate the referred and related objects in the prompt. For example, the query “car to the left of a pedestrian”, has high similarity with both “the vehicle is to the left of the person” and “the person is to the right of the vehicle”. To resolve this issue, we post-process predictions with Filter by Referred Class to obtain the final set of referred objects.
@@
Referential Tracking by Program Synthesis. RefProg filters off-the-shelf tracks using Python programs synthesized by a large language model (LLM), without explicitly relying on any information from LiDAR or cameras. To reduce the time required to run these synthesized programs, we first filter the input tracks based on confidence. We compute each track’s confidence score by summing its confidence values across all timestamps. For the classes REGULAR_VEHICLE, PEDESTRIAN, BOLLARD, CONSTRUCTION_CONE, and CONSTRUCTION_BARREL, we retain the top 200 tracks per class. For all other classes, we retain the top 100 tracks per category. We prompt the LLM as follows:
@@
Please use the following functions to find instances of a referred object in an autonomous driving dataset. Be precise to the description, try to avoid returning false positives. API Listing: {API Listing} Categories: {AV2 Categories} Define a single scenario for the description:{Natural Language Description} Here is a list of examples: {Prediction Examples}. Only output code and comments as part of a Python block. Do not define any additional functions, or filepaths. Do not include imports. Assume the log_dir and output_dir variables are given. Wrap all code in one python block and do not provide alternatives. Output code even if the given functions are not expressive enough to find the scenario.
@@
We include the API listing in Appendix J. We found this prompt consistently produces reasonable code across all tested LLMs. After running the synthesized programs, we post-process the output tracks. We discard object relationships where the spatial distance exceeds 50 meters. For the remaining tracks, we apply timestamp dilation to reduce flickering of the referred object labels. We symmetrically dilate time segments to a minimum of 1.5 seconds. For example, if the base tracker tracks an object with ID 1 from 3.0 to 15.0 seconds at 2 Hz, and the synthesized code marks it as referred at timestamps 4.5 and 5.0, we update the final output to label the track as referred from 4.0 to 5.5 seconds. Finally, we downsample all output tracks to 2 Hz for evaluation.
Appendix B Token Statistics for LLM Baselines
ReferGPT, LLM APIs as a Black Box, and RefProg can be viewed as different ways of providing scaffolding around an LLM for RMOT and scenario mining tasks. Most commercially available LLM APIs bill according to four cost categories: uncached input tokens, cached input tokens, output tokens, and code interpreter sessions (CIS). Input token count scales approximately linearly with the length of the input text, the number of input images, and seconds of input video. ReferGPT requires orders of magnitude more input and output tokens because it calls a VLM call for every tracked object at every timestamp. It also requires processing images, which can use more than 1000 tokens (depending on resolution) using GPT-5-mini. LLM APIs as a Black Box only makes as many calls as are there logs of data. However, inspecting the reasoning traces reveals that GPT-5 automatically went through multiple rounds of generating data processing code and verifying the output. The LLM also often re-generates the same processing code between logs. Initializing code interpreter sessions represents approximately half of the total cost ($) of inference. RefProg makes as many calls as there are unique scenario descriptions. Since the entire atomic function API must be given in context, it requires a relatively high number of input tokens. However, most of these tokens are cached between requests. The output of the RefProg is relatively compact because the LLM has access to high-level composable functions and does not need to write extensive Python code.
| Method | LLM | HOTA-Temporal | Cost | Uncached Input | Cached Input | Output | CIS |
| ReferGPT | GPT-5-mini | 20.7 | 150.64 | 372.9M | 4.0M | 28.7M | 0 |
| LLM API as Black Box | GPT-5 | 37.2 | 77.59 | 4.5M | 0 | 3.4M | 1283 |
| RefProg (Ours) | GPT-5 | 42.3 | 13.89 | 0.6M | 2.6M | 1.3M | 0 |
Appendix C Comparison of Recent Datasets
We provide a qualitative comparison of language expressions found in RefAV with other popular RMOT-for-driving and VQA-for-driving datasets. In particular, we examine nuPrompt, Refer-KITTI, and nuScenes-QA. To obtain a representative sample of each dataset, we compute text embeddings for each expression using Qwen-3-Embedding-8B. We display the first 30 samples using farthest point sampling initialized from the two samples furthest apart from each other. Many expressions in nuPrompt deal with object attributes that are minimally relevant to driving such as car color, a person’s gender, or a person’s clothing. Refer-KITTI similarly only refers to annotated vehicle and pedestrian classes. nuScenes-QA asks questions that are relevant to driving, but the dataset is limited by its lack of human annotations and unnatural sounding questions. We present examples from each dataset below.
RefAV (Ours)
nuPrompt [wu2023nuprompt]
Refer-KITTI [wu2023referkitti]
nuScenes-QA [qian2024nuscenesqa]
Appendix D Scenario Mining Annotation Tool